An office have 2 types of employees

administrators (female=25, male=15)
operators (female=35, male=25)

If one person chosen at random, what is the probability that the person is

1) a male administrator
2) a female operator
3) male
4) an operator
1
Expert's answer
2018-03-01T09:17:08-0500
1) The probability that the person is a male administrator: P = 15/(25+15+35+25)= 15/100 = 0.15.
2) The probability that the person is a female operator: P =35/(25+15+35+25)= 35/100 = 0.35.
3) The probability that the person is male: P =(15+25)/(25+15+35+25)= 40/100 = 0.4.
4) The probability that the person is an operator: P =(35+25)/(25+15+35+25)= 60/100 = 0.6.